A non-relativistic particle of mass m and charge e , moving with a velocity v
and
acceleration a
,emits radiation of intensity I . What is the intensity of the radiation
emitted by a particle of mass m / 2 , charge 2e , velocity v / 2 and acceleration 2a?
1
Expert's answer
2020-09-28T08:09:31-0400

Solution

According to given data of question intensity can be expressed as below

"I=A.\frac{q^2 a^2 sin^2\theta}{r^2}"

Then intensities ratio can be expressed as

"\frac{I_2}{I_1}=\frac{q_2^2 a_2^2 }{q_1^2 a_1^2}"

Using the all given data in question

Intensity in second case is written as

"{I_2}=I_1\frac{4e^2 4a^2 }{e^2 a^2}"

"I_2=16I" where I1=I

Therefore intensity is 16I in second case.